### 12 Apr 2026 v4

# LPNSR: Optimal Noise-Guided Diffusion Image Super-Resolution Via Learnable Noise Prediction

Abstract
Diffusion-based image super-resolution (SR) aims to reconstruct high-resolution (HR) images from low-resolution (LR) observations, yet faces a fundamental trade-off between inference efficiency and reconstruction quality in limited-step sampling scenarios. A critical yet underexplored question is: what is the optimal noise to inject at each intermediate diffusion step? In this paper, we establish a theoretical framework that derives the closed-form analytical solution for optimal intermediate noise in diffusion models from a maximum likelihood estimation perspective, revealing a consistent conditional dependence structure that generalizes across diffusion paradigms. We instantiate this framework under the residual-shifting diffusion paradigm and accordingly design an LR-guided multi-input-aware noise predictor to replace random Gaussian noise. We further mitigate initialization bias with a high-quality pre-upsampling network. The compact 4-step trajectory uniquely enables end-to-end optimization of the entire reverse chain, which is computationally prohibitive for conventional long-trajectory diffusion models. Extensive experiments demonstrate that LPNSR achieves state-of-the-art perceptual performance on both synthetic and real-world datasets, without relying on any large-scale text-to-image priors. The source code of our method can be found at https://github.com/Faze-Hsw/LPNSR.
1 Introduction
Image super-resolution (SR) aims to recover high-resolution (HR) images from low-resolution (LR) observations, a severely ill-posed problem due to unknown real-world degradations. Recently, diffusion models(Saharia et al., 2023; Yue et al., 2023; Wang et al., 2024a; Yue et al., 2025; Kawar et al., 2022; Chung et al., 2022c; Rombach et al., 2022; Wu et al., 2024a, b; Ho et al., 2020) have demonstrated unprecedented potential in SR tasks, achieving remarkable breakthroughs in both pixel-level fidelity and perceptual realism. However, diffusion-based SR methods face a fundamental and critical trade-off between inference efficiency and reconstruction performance, especially in limited-step sampling scenarios that are essential for practical deployment.
To break this trade-off, the residual-shifting diffusion framework (ResShift(Yue et al., 2023)) has emerged as the state-of-the-art (SOTA) efficient solution, achieving SR inference with only 4 sampling steps while retaining a lightweight denoising network. However, due to sampling step compression, its 4-step version suffers from severe performance degradation compared to the 15-step counterpart. This performance decline exposes a fundamental flaw in mainstream diffusion pipelines(Yue et al., 2023; Lu et al., 2022; Nichol and Dhariwal, 2021; Song et al., 2020; Ho et al., 2020): the universal use of unconstrained random Gaussian noise in intermediate reverse steps. Existing attempts to address this issue mainly rely on distillation to compress the diffusion sampling trajectory into a single step(Wang et al., 2024b; Wu et al., 2024a), which avoids introducing random Gaussian noise in intermediate steps, yet the performance is inherently bounded by the capacity of the teacher model. Diffusion inversion methods(Chung et al., 2022b, a; Fei et al., 2023; Song et al., 2023; Xiao et al., 2024; Yue and Loy, 2024) adopt step-wise optimization for intermediate noise. However, existing diffusion inversion methods mainly focus on image editing and lack a closed-form definition of optimal noise with a generalizable structure across diffusion paradigms. In this work, we address these gaps by establishing a unified maximum likelihood estimation (MLE)-based theoretical framework to derive the closed-form analytical solution of optimal intermediate noise for diffusion models, revealing its consistent conditional dependence structure generalizable across mainstream diffusion paradigms, laying a theoretical foundation for mitigating few-step diffusion performance degradation.
We instantiate this framework on the residual-shifting diffusion paradigm, as its compact 4-step trajectory uniquely enables end-to-end optimization of the full reverse chain—computationally infeasible for long-trajectory models like DDPM(Ho et al., 2020). This allows us to learn a deep neural network for the theoretically optimal noise without modifying the pretrained denoising network or breaking the original efficient residual-shifting mechanism. We further address initialization bias from bicubic upsampling, a critical bottleneck for few-step sampling where limited iterations cannot correct initial deviations, via a pretrained pre-upsampling network to generate a high-quality initial state before diffusion. An additional key benefit of this design is that it enables arbitrary-step super-resolution inference without redesigning the diffusion framework’s hyper parameters or retraining the pretrained denoising network.
Building on this, we propose LPNSR, an efficient prior-enhanced diffusion SR framework. It adopts an LR-guided multi-input-aware noise predictor (aligned with our derived optimal noise structure) to replace random Gaussian noise, paired with the pre-upsampling initialization strategy. Extensive experiments show LPNSR achieves state-of-the-art perceptual SR performance without external text-to-image (T2I) priors.
The main contributions of this work are as follows:
• We establish a unified MLE-based framework to derive the closed-form optimal intermediate noise for diffusion models, revealing its generalizable conditional dependence structure across paradigms.
• We instantiate this framework on the residual-shifting diffusion paradigm, designing an LR-guided noise predictor to approximate the optimal noise while fully preserving the original efficient inference mechanism.
• We mitigate few-step initialization bias via a pretrained pre-upsampling network, which significantly boosts compact-trajectory inference performance and enables flexible arbitrary-step inference without retraining the core denoising network.
Figure 1: Qualitative comparison of our PreSet-A and PreSet-B methods under different sampling steps for $\times 4$ image super-resolution. (a) Zoomed patch of the input LR image; (b)-(e) Results of PreSet-A with 4, 3, 2, and 1 sampling steps, respectively; (f)-(i) Results of PreSet-B with 4, 3, 2, and 1 sampling steps, respectively. Two representative samples are provided to demonstrate the visual performance of different configurations. (Zoom in for best view)
2 Related work
Image Super-Resolution. Along with the proliferation of deep learning, deep learning-driven approaches have progressively emerged as the dominant paradigm for SR(Dong et al., 2015; Rojas Sedó, 2022). Early prominent works primarily focused on training regression models using paired LR-HR data(Ahn et al., 2018; Kim et al., 2016; Wang et al., 2015). Though these models effectively capture the expectation of the posterior distribution, they inherently suffer from over-smoothing artifacts in generated results(Ledig et al., 2017; Menon et al., 2020; Sajjadi et al., 2017). To enhance the perceptual quality of reconstructed HR images, generative SR models have garnered growing interest—including autoregressive architectures(Dahl et al., 2017; Menick and Kalchbrenner, 2018; Van den Oord et al., 2016; Parmar et al., 2018). Despite notable gains in perceptual performance, autoregressive models typically incur substantial computational overhead. Additionally, GAN-based SR methods have attained remarkable success in perceptual quality(Guo et al., 2022; Karras et al., 2017; Ledig et al., 2017; Menon et al., 2020; Sajjadi et al., 2017), yet the training process remains unstable. More recently, diffusion-based models have become a focal point of SR research(Choi et al., 2021; Chung et al., 2022c; Kawar et al., 2022; Rombach et al., 2022; Saharia et al., 2023). These methods generally fall into two categories: those that concatenate the LR image to the denoiser’s input(Rombach et al., 2022; Saharia et al., 2023), and those that adapt the backward process of a pre-trained diffusion model(Choi et al., 2021; Chung et al., 2022c; Kawar et al., 2022). While these diffusion-based approaches yield promising performance, their methods still introduce unconstrained random Gaussian noise in each step of the reverse diffusion process, rather than meaningful noise maps.
Diffusion Inversion. This paradigm aims to find the optimal noise maps that reconstruct the target image when fed into a diffusion model. Early works optimized text embeddings for better alignment(Gal et al., 2022; Mokady et al., 2023), and follow-up works further refined inversion via textual/visual prompts(Miyake et al., 2025; Nguyen et al., 2023) or intermediate noise map optimization(Ju et al., 2023; Kang et al., 2024; Meiri et al., 2023; Wallace et al., 2023). However, existing diffusion inversion methods are mostly heuristic step-wise optimization schemes tailored for image editing tasks. None of them establish a unified, generalizable theoretical derivation paradigm for optimal intermediate noise, nor provide a closed-form analytical solution that can generalize across mainstream diffusion paradigms. For SR tasks, InvSR(Yue et al., 2025) extended diffusion inversion to SR, but it is constrained by the long sampling trajectory of DDPM(Ho et al., 2020), which makes full-chain end-to-end optimization computationally prohibitive. It only optimizes the initial noise map, and fails to solve the fundamental problem of defining and optimizing the optimal intermediate noise for the full reverse sampling chain.
Figure 2: Visualization of the intermediate noise maps generated by our proposed noise predictor during the 4-step reverse diffusion process. From left to right: the input LR image, and the predicted noise maps at step-4, step-3, and step-2 of the reverse sampling process, respectively.
Figure 3: Statistical distribution analysis of the outputs from our LR-guided noise predictor. From left to right: the input LR image, the final SR image generated by LPNSR, the probability density distributions of the predicted noise maps at each intermediate reverse step ($t=4$, $t=3$, and $t=2$), and the distribution of the final SR output in latent space. The mean ($\mu$) and standard deviation ($\sigma$) of the noise/latent values are provided for each distribution.
3 Methodology
3.1 Peliminaries
We first establish a unified notation system for conditional diffusion models, which is generalizable to mainstream diffusion paradigms including DDPMHo et al. (2020) and residual-shifting frameworkYue et al. (2023), to lay a formal foundation for our theoretical derivation. We denote $x_{0}$ as the clean target sample (i.e., HR image in the SR task), and $y$ as the task conditional input (i.e., the LR image in the SR task). The forward diffusion process is defined as a Markov chain of length $T$, which gradually corrupts the clean sample $x_{0}$ with noise. The single-step transition distribution follows an isotropic Gaussian distribution:
$q(x_{t}|x_{t-1},y)=\mathcal{N}\left(x_{t};\boldsymbol{\mu}_{t}^{f}(x_{t-1},y),\Sigma_{t}^{f}\cdot\boldsymbol{I}\right),\quad t=1,2,...,T,$ (1)
where $\boldsymbol{\mu}_{t}^{f}(\cdot)$ and $\Sigma_{t}^{f}$ are the forward transition mean function and scalar variance, specified by the target diffusion paradigm. For most mainstream paradigms, the forward process admits a closed-form marginal distribution, which directly gives the distribution of the noisy sample at arbitrary timestep t without iterative sampling:
$q(x_{t}|x_{0},y)=\mathcal{N}\left(x_{t};\boldsymbol{\mu}_{t}^{m}(x_{0},y),\Sigma_{t}^{m}\cdot\boldsymbol{I}\right),\quad t=1,2,...,T,$ (2)
where $\boldsymbol{\mu}_{t}^{m}(\cdot)$ and $\Sigma_{t}^{m}$ denoting the marginal mean function and scalar variance, respectively. The learnable reverse denoising process is an inverse Markov chain that reconstructs $x_{t-1}$ from $x_{t}$ under the guidance of $y$, with Gaussian distribution:
$p_{\theta}(x_{t-1}|x_{t},y)=\mathcal{N}\left(x_{t-1};\boldsymbol{\mu}_{t}^{\theta}(x_{t},y,t),\Sigma_{t}\cdot\boldsymbol{I}\right),$ (3)
where $\boldsymbol{\mu}_{t}^{\theta}(\cdot)$ is the reverse mean parameterized by a learnable denoising network $f_{\theta}$, and $\Sigma_{t}$ is the fixed or predefined reverse scalar variance. This gives the unified single-step reverse iteration:
$x_{t-1}=\boldsymbol{\mu}_{t}^{\theta}(x_{t},y,t)+\sqrt{\Sigma_{t}}\cdot z_{t-1},$ (4)
where $z_{t-1}$ is the intermediate noise injected at step $t$, universally sampled from an unconstrained standard Gaussian distribution $\mathcal{N}(\mathbf{0},\mathbf{I})$ in conventional pipelines, which is the core variable we optimize in this work.
3.2 Theoretical derivation of optimal intermediate noise
We define the optimal intermediate noise $z_{t-1}^{*}$ as the noise that maximizes the conditional log-likelihood of the ground-truth $x_{0}$ given the intermediate state $x_{t}$ and conditional input $y$, following the MLE paradigm for generative model optimization:
$z_{t-1}^{*}=\arg\max_{z_{t-1}}\log p_{\theta}\left(x_{0}|x_{t-1}(z_{t-1}),y\right),$ (5)
where $x_{t-1}(z_{t-1})$ is uniquely determined by $z_{t-1}$ via Eq. (4). By Bayes’ theorem, we decompose the posterior as
$p_{\theta}\left(x_{0}|x_{t-1},y\right)=\frac{q\left(x_{t-1}|x_{0},y\right)\cdot p_{\theta}\left(x_{0}|y\right)}{p_{\theta}\left(x_{t-1}|y\right)}.$ (6)
Following standard diffusion posterior derivation practices(Ma et al., 2023), we adopt a non-informative prior $p_{\theta}\left(x_{0}|y\right)\propto 1$, and the marginal likelihood $p_{\theta}\left(x_{t-1}|y\right)$ is a normalization constant independent of $x_{0}$. This simplifies the posterior to $p_{\theta}\left(x_{0}|x_{t-1},y\right)\propto q\left(x_{t-1}|x_{0},y\right)$. Substituting the forward marginal Gaussian distribution in Eq. (2) and taking the logarithm, the log-likelihood becomes
$\log p_{\theta}\left(x_{0}|x_{t-1},y\right)\propto-\frac{1}{2\Sigma_{t-1}^{m}}\left\|x_{0}-\left(\boldsymbol{\mu}_{t-1}^{m}\right)^{-1}\left(x_{t-1}\right)\right\|^{2},$ (7)
where $\left(\boldsymbol{\mu}_{t-1}^{m}\right)^{-1}\left(\cdot\right)$ is the inverse mapping of the marginal mean function with respect to $x_{0}$. Maximizing the log-likelihood is equivalent to minimizing the above L2 norm, leading to the optimization objective:
$z_{t-1}^{*}=\arg\min_{z_{t-1}}\left\|x_{0}-\left(\boldsymbol{\mu}_{t-1}^{m}\right)^{-1}\left(x_{t-1}(z_{t-1})\right)\right\|^{2}.$ (8)
Substituting Eq. (4) into this objective and solving the convex optimization problem, we obtain the closed-form analytical solution of the optimal intermediate noise for general conditional diffusion models:
$z_{t-1}^{*}=\frac{\boldsymbol{\mu}_{t-1}^{m}(x_{0},y)-\boldsymbol{\mu}_{t}^{\theta}(x_{t},y,t)}{\sqrt{\Sigma_{t}}}.$ (9)
This closed-form solution yields three fundamental conclusions. First, the optimal intermediate noise is a deterministic mapping rather than an unconstrained random Gaussian variable, directly proving the inherent suboptimality of conventional random noise injection, especially in compact few-step sampling scenarios with severe error accumulation. Second, the optimal noise follows a unified conditional dependence structure across all mainstream diffusion paradigms, uniquely determined by the forward marginal mean $\boldsymbol{\mu}_{t-1}^{m}(x_{0},y)$, reverse mean $\boldsymbol{\mu}_{t}^{\theta}(x_{t},y,t)$, and reverse variance $\sqrt{\Sigma_{t}}$. Third, the solution explicitly defines the input variables required to approximate the optimal noise, providing theoretical guidance for the design of our noise prediction network.
3.3 Instantiation to residual-shifting diffusion SR paradigm
We instantiate our general optimal noise derivation paradigm proposed in Sec. 3.2 to the residual-shifting diffusion frameworkYue et al. (2023), a efficient SR pipeline with a compact 4-step sampling trajectory. Unlike long-trajectory models (e.g., DDPMHo et al. (2020)) requiring hundreds of steps, this 4-step design uniquely enables end-to-end optimization of the full reverse chain, which is computationally prohibitive for conventional long-trajectory diffusion models.
Optimal Intermediate Noise. We follow the native notation of the residual-shifting paradigm: $y_{0}$ is the input LR image, $x_{0}$ is the target HR image, and $e_{0}=y_{0}-x_{0}$ is the LR-HR residual. It’s assumed that the input LR image $y_{0}$ shares the identical spatial dimension with the target HR image $x_{0}$, which can be achieved by upsampling the raw LR input. Its forward marginal distribution are
$q(x_{t}|x_{0},y_{0})=\mathcal{N}\left(x_{t};\underbrace{(1-\eta_{t})x_{0}+\eta_{t}y_{0}}_{\boldsymbol{\mu}_{t}^{m}(x_{0},y_{0})},\underbrace{\kappa^{2}\eta_{t}}_{\Sigma_{t}^{m}}\cdot\boldsymbol{I}\right),$ (10)
where $\{\eta_{t}\}_{t=1}^{T}$ is a monotonically increasing shifting sequence ($\eta_{1}\to 0$, $\eta_{T}\to 1$), $\kappa$ is the noise variance hyperparameter. And its reverse process can be given as
$p_{\theta}(x_{t-1}|x_{t},y_{0})=\mathcal{N}\left(x_{t-1};\underbrace{\frac{\eta_{t-1}}{\eta_{t}}x_{t}+\frac{\alpha_{t}}{\eta_{t}}f_{\theta}(x_{t},y_{0},t)}_{\boldsymbol{\mu}_{t}^{\theta}(x_{t},y_{0},t)},\underbrace{\kappa^{2}\frac{\eta_{t-1}\alpha_{t}}{\eta_{t}}}_{\Sigma_{t}}\cdot\boldsymbol{I}\right),$ (11)
where $\alpha_{t}=\eta_{t}-\eta_{t-1}$ ($\alpha_{1}=\eta_{1}$), $f_{\theta}$ is the pretrained denoiser predicting clean image $x_{0}$. Substituting the above marginal mean, reverse mean and variance into our general optimal noise solution Eq. (9), we directly obtain the closed-form optimal noise for this paradigm:
$z_{t-1}^{*}=\frac{(1-\eta_{t-1})x_{0}+\eta_{t-1}y_{0}-\boldsymbol{\mu}_{t}^{\theta}(x_{t},y_{0},t)}{\sqrt{\Sigma_{t}}}.$ (12)
This solution proves the inherent suboptimality of the original random Gaussian noise injection, and explicitly defines that the optimal noise is uniquely determined by $\{x_{t},x^{\prime}_{0}=f_{\theta}(x_{t},y_{0},t),y_{0},t\}$, providing theoretical guidance for our noise predictor design.
Initialization Strategy. For the residual-shifting paradigm, the forward marginal distribution at $t=T$ converges to $q(x_{T}|x_{0},y_{0})=\mathcal{N}(y_{0},\kappa^{2}\boldsymbol{I})$. During inference, the ground-truth $x_{0}$ is inaccessible, so we directly replace $x_{0}$ in the forward marginal distribution $q(x_{t}|x_{0},y_{0})$ with $y_{0}$, yielding the arbitrary-step initialization formula:
$x_{t}=y_{0}+\kappa\sqrt{\eta_{t}}\cdot z_{t},z_{t}\sim\mathcal{N}(\boldsymbol{0},\boldsymbol{I}).$ (13)
The validity of this approximation hinges on the proximity between $y_{0}$ and $x_{0}$. The native pipeline uses naive bicubic upsampling to match $y_{0}$ to $x_{0}$’s dimension, but this fails to bring $y_{0}$ sufficiently close to ground-truth HR image $x_{0}$, introducing severe initialization bias. In compact few-step trajectories, the model lacks enough iterations to correct this bias, causing dramatic performance degradation. We thus replace bicubic upsampling with a pretrained SwinIR-GAN(Liang et al., 2021) to generate dimension-matched $y_{0}$, narrowing the $y_{0}-x_{0}$ gap, mitigating initialization bias, boosting few-step performance, and enabling 1-4 step arbitrary inference without retraining the core denoiser (see Tab. 1).
End-to-End Optimization. Since $x_{0}$ is unavailable during inference, we adopt the UNetRonneberger et al. (2015) used in Yue et al. (2023) as the noise predictor $g_{\omega}$ to approximate the optimal noise in Eq. (12), taking $\{x_{t},x^{\prime}_{0},y_{0},t\}$ as input. The revised reverse iteration is
$x_{t-1}=\boldsymbol{\mu}_{t}^{\theta}(x_{t},y_{0},t)+\sqrt{\Sigma_{t}}\cdot g_{\omega}(x_{t},x_{0}^{\prime},y_{0},t).$ (14)
Substituting the optimal noise into the reverse process (Eq. (11)) gives $x_{t-1}$ that exactly matches the conditional mean of the forward marginal distribution (Eq. (10)), meaning injecting optimal noise at every step enforces the reverse trajectory to align with the forward process, and guarantees exact recovery of the ground-truth HR image for a well-trained denoiser $f_{\theta}$. To enable $g_{\omega}$ to learn this optimal mapping, we optimize it end-to-end over the full 4-step reverse chain (computationally feasible only for this compact trajectory), freezing the pretrained denoiser $f_{\theta}$ and the VQGAN autoencoderEsser et al. (2021). Following recent SR approachesSaharia et al. (2023); Yue et al. (2025); Wang et al. (2021), the training objective is a combination of L1 loss $L_{1}$, LPIPSZhang et al. (2018) loss $L_{l}$, and GANGoodfellow et al. (2014) loss $L_{g}$:
$\mathcal{L}=\lambda_{1}L_{1}(x_{0}^{\prime},x_{0})+\lambda_{l}L_{l}(x_{0}^{\prime},x_{0})+\lambda_{g}L_{g}(x_{0}^{\prime},x_{0}),$ (15)
where $x_{0}^{\prime}$ is the final predicted clean image. $\lambda_{1}$, $\lambda_{l}$, and $\lambda_{g}$ are hyperparameters balancing the contributions of each loss component. The detailed training and inference procedure is provided in Alg. 1 and Alg. 2.
Figure 4: Visual results of different methods on two typical real-world examples. (Zoom in for best view)
Table 1: Quantitative comparison results between our proposed methods (denoted as PreSet-A, PreSet-B) and the original ResShift on the ImageNet-Test dataset, where PreSet-A uses only the noise predictor, and PreSet-B employs SwinIR-GAN to do pre-upsampling. The Runtime metric denotes the average inference time per image, which is tested on a single NVIDIA RTX 3090 Ti GPU. (Notably, the noise predictor is not activated during single-step inference, thus PreSet-A yields identical inference results to ResShift.)
Steps Methods Metrics
[…]
$T=1$ ResShift 28.96 0.7886 0.4183 7.7442 7.1397 0.2929 30.2567 0.64
PreSet-A 28.96 0.7886 0.4183 7.7442 7.1397 0.2929 30.2567 0.64
PreSet-B 27.11 0.7566 0.2185 5.2409 4.2176 0.5776 66.0646 0.81
$T=2$ ResShift 28.48 0.7823 0.3335 7.1880 6.7002 0.3392 38.8335 0.69
PreSet-A 28.01 0.7629 0.2861 5.9066 5.7823 0.3812 44.2488 0.71
PreSet-B 25.94 0.7244 0.2214 4.2463 3.2883 0.6341 70.7036 0.89
$T=3$ ResShift 28.62 0.7816 0.2487 6.1413 5.7854 0.4601 52.4232 0.74
PreSet-A 26.68 0.7065 0.2561 4.4798 3.2974 0.6557 66.7649 0.77
PreSet-B 25.99 0.7001 0.2575 4.4535 3.2303 0.6858 71.5964 0.98
$T=4$ ResShift 27.33 0.7530 0.1998 5.8700 4.3643 0.6147 65.5860 0.81
PreSet-A 26.35 0.7151 0.2324 4.4127 3.2834 0.6689 71.2560 0.90
PreSet-B 26.11 0.7054 0.2424 4.3807 3.1995 0.6921 71.7105 1.09
Table 2: Quantitative comparisons of different methods on ImageNet-Test and RealSR datasets. The best and second-best results are highlighted in red and blue.
Metrics
[…]
ImageNet-Test BSRGANZhang et al. (2021) 27.05 0.7453 0.2437 4.5345 3.7111 0.5703 67.7195
RealESRGANWang et al. (2021) 26.62 0.7523 0.2303 4.4909 3.7234 0.5090 64.8186
SeeSRWu et al. (2024b) 26.69 0.7422 0.2187 4.3825 3.4742 0.5868 71.2412
ResShiftYue et al. (2023) 27.33 0.7530 0.1998 5.8700 4.3643 0.6147 65.5860
SinSRWang et al. (2024b) 26.98 0.7304 0.2209 5.2623 3.8189 0.6618 67.7593
OSEDiffWu et al. (2024a) 23.95 0.6756 0.2624 4.7157 3.3775 0.6818 70.3928
InvSRYue et al. (2025) 24.14 0.6789 0.2517 4.3815 3.0866 0.7093 72.2900
LPNSR(Ours) 26.11 0.7054 0.2424 4.3807 3.1995 0.6921 71.7105
RealSR BSRGANZhang et al. (2021) 26.51 0.7746 0.2685 4.6501 4.4644 0.5439 63.5869
RealESRGANWang et al. (2021) 25.85 0.7734 0.2728 4.6766 4.4881 0.4898 59.6803
SeeSRWu et al. (2024b) 26.20 0.7555 0.2806 4.5358 4.1464 0.6824 66.3757
ResShiftYue et al. (2023) 25.77 0.7453 0.3395 6.9113 5.4013 0.5994 57.5536
SinSRWang et al. (2024b) 26.02 0.7097 0.3993 6.2547 4.7183 0.6634 59.2981
OSEDiffWu et al. (2024a) 23.89 0.7030 0.3288 5.3310 4.3584 0.7008 65.4806
InvSRYue et al. (2025) 24.50 0.7262 0.2872 4.2189 3.7779 0.6918 67.4586
LPNSR(Ours) 24.62 0.7003 0.3229 4.2175 3.6963 0.7180 67.5634
Table 3: Quantitative comparisons of various methods on RealSet80 dataset. The best and second-best results are highlighted in red and blue.
[…]
BSRGANZhang et al. (2021) 4.4408 4.0276 0.6263 66.6288
RealESRGANWang et al. (2021) 4.1568 3.8852 0.6189 64.4957
SeeSRWu et al. (2024b) 4.3678 3.7429 0.7114 69.7658
ResShiftYue et al. (2023) 5.9866 4.8318 0.6515 61.7967
SinSRWang et al. (2024b) 5.6243 4.2830 0.7228 64.0573
OSEDiffWu et al. (2024a) 4.3457 3.8219 0.7093 68.8202
InvSRYue et al. (2025) 4.0284 3.4666 0.7291 69.8055
LPNSR(Ours) 4.3066 3.5845 0.7316 70.2184
Table 4: Ablation study results of noise predictor at each intermediate step on the RealSR dataset. We evaluate the performance of LPNSR when replacing the noise predictor with random Gaussian noise at $t=4$, $t=3$, and $t=2$ individually, under the 4-step sampling setting.
[…]
LPNSR w/o Predictor at $t=4$ 24.53 0.6898 0.3434 4.3864 3.7770 0.7090 66.7373
LPNSR w/o Predictor at $t=3$ 24.05 0.6848 0.3404 4.3860 3.7553 0.7374 67.3563
LPNSR w/o Predictor at $t=2$ 24.86 0.7308 0.3117 5.8530 4.7332 0.7041 63.5838
LPNSR 24.62 0.7003 0.3229 4.2175 3.6963 0.7180 67.5634
4 Experiments
In this section, we compare our method against some of the recent state-of-the-art diffusion-based SR approaches, analyze the effectiveness of our LR-guided noise predictor, and perform ablation studies to understand the contributions of different components in our model. Our experiments mainly focus on the $\times 4$ SR task.
4.1 Experimental setup
Training Details. We train the noise predictor on the LSDIRLi et al. (2023) dataset and the first 10k face images from the FFHQKarras et al. (2019) dataset for over 200k iterations, randomly cropping an image patch with a resolution of $256\times 256$ from the source image and synthesizing the LR image using the pipeline of RealESRGANWang et al. (2021) at each iteration. We adopt the AdamWLoshchilov and Hutter (2017) optimizer with a learning rate of $5\times 10^{-5}$ and a batch size of 16, while using the CosineAnnealingLoshchilov and Hutter (2016) scheduler with a minimum learning rate of $1\times 10^{-5}$. The hyperparameters for the loss function are set as $\lambda_{1}=1.0$, $\lambda_{l}=1.0$, and $\lambda_{g}=0.1$. During training, we set $T=4$ to remain consistent with ResShiftYue et al. (2023),and the noise variance hyperparameter $\kappa=2.0$ as well as the shifting sequence $\{\eta_{t}\}_{t=1}^{T}$ also follow the identical settings. The denoising network $f_{\theta}$ and the VQGAN autoencoderEsser et al. (2021) is frozen during training, only the noise predictor is optimized.
Testing Datasets and Metrics. To facilitate fair and direct comparison with the latest SOTA methods, we follow the experimental setup of InvSRYue et al. (2025) by adopting its testing datasets and evaluation metrics. Specifically, our experiments are conducted on the three datasets: the synthetic dataset ImageNet-TestDeng et al. (2009) used in Yue et al. (2025), the real-world datasets RealSRCai et al. (2019) and RealSet80Yue et al. (2023). For evaluation metrics, we retain the same configuration:seven metrics (three reference metrics: PSNR, SSIMWang et al. (2004), LPIPSZhang et al. (2018); four non-reference metrics: NIQEMittal et al. (2012), PIBlau et al. (2018), MUSIQKe et al. (2021), CLIPIQAWang et al. (2023)) are employed for ImageNet-Test and RealSR,while only non-reference metrics are used for RealSet80. PSNR and SSIM are calculated on the luminance (Y) channel of YCbCr space, and other metrics are computed in the standard sRGB space.
Compared Methods. To benchmark our model, we compare it against eight recent methods: 2 GAN-based methods (BSRGANZhang et al. (2021), RealESRGANWang et al. (2021)) and 5 diffusion-based methods (SeeSRWu et al. (2024b), ResShiftYue et al. (2023), SinSRWang et al. (2024b), OSEDiffWu et al. (2024a), InvSRYue et al. (2025)). The presets of all methods follow the official default guidelines.
4.2 Experimental results
Inference Steps. Tab. 1 compares our PreSet-A (noise predictor only) and PreSet-B (with pre-upsampling) methods against original ResShiftYue et al. (2023) across 1-4 sampling steps. We observe that the pre-upsampling module delivers more significant gains with fewer steps, while our noise predictor stably improves perceptual performance across all settings, with all methods performing better with more steps. Regarding inference latency, our noise predictor introduces no noticeable overhead as its computation is performed entirely in the latent space. In contrast, the pre-upsampling network operates in the image space and thus incurs additional runtime. For 3-4 step inference, the model has sufficient iterations to mitigate initialization bias, so the pre-upsampling module can be optionally omitted to trade off a minor performance drop for faster inference. For 1-2 step inference, however, the pre-upsampling module is indispensable, as the limited sampling steps cannot compensate for initialization deviation without high-quality pre-upsampling. Qualitative results in Fig. 1 confirm that PreSet-B maintains consistent visual quality even in 1-2 step inference, while PreSet-A suffers from blurriness in low-step settings. We fix 4-step inference with pre-upsampling as our default LPNSR configuration for all following experiments.
Performance Comparison. Tab. 2 and 3 presents a comprehensive comparison of our LPNSR against recent SOTA methods on the ImageNet-Test, RealSR and RealSet80 datasets. Compared to the baseline ResShiftYue et al. (2023), our LPNSR achieves remarkable improvements in perceptual metrics (e.g., NIQE, CLIPIQA, MUSIQ) while maintaining competitive fidelity. Against T2I-utilizing models such as OSEDiffWu et al. (2024a), InvSRYue et al. (2025) and SeeSRWu et al. (2024b), LPNSR delivers comparable or better perceptual quality without leveraging any pre-trained text-to-image priors. It also outperforms multi-step diffusion methods (e.g.,SeeSRWu et al. (2024b)) on core perceptual metrics. On real-world datasets, LPNSR ranks among the top-tier SOTA methods. It achieves leading perception-oriented metrics, such as NIQE, PI, CLIPIQA and MUSIQ on RealSR. On RealSet80, LPNSR attains top-2 PI score, the best MUSIQ and CLIPIQA scores among all competing methods. Qualitatively, Fig. 4 shows LPNSR generates sharper textures and more consistent structures than other methods, free from spurious details or over-smoothing (see Appendix for more visual comparisons). LPNSR generates SR images with sharp details, intact structural consistency, and no noticeable artifacts. It effectively restores natural textures and clear edge contours that align with the input LR structure, delivering visually coherent and realistic results.
Analysis and Ablation Study of Noise Predictor. Across the 4-step coarse-to-fine denoising trajectory, our LR-guided noise predictor implements progressive prior guidance aligned with the denoising logic. As shown in Fig. 2, the predicted noise maps are highly aligned with the LR image’s structure and texture, following a hierarchical guidance pattern: the step 4 map anchors global structure to avoid initial sampling deviation, the step 3 map focuses on mid-frequency texture refinement to suppress cumulative error, and the step 2 map targets local fine-grained details to optimize perceptual quality. Statistical distribution analysis in Fig. 3 further validates this progressive, LR-aligned variation pattern, and confirms that the optimized noise distributions at all steps are not unconstrained random Gaussian distributions, verifying the intrinsic working mechanism of our predictor. We further conduct a step-wise ablation study on the RealSR dataset (see Tab. 4) to quantitatively verify the independent contribution of each step’s predictor: the full LPNSR model achieves the best overall performance, validating the effectiveness of our full-stage prior guidance; removing the step 4 predictor degrades both fidelity and perceptual quality, disabling the step 3 predictor causes the most severe PSNR drop, and replacing the step 2 predictor with random noise sharply impairs perceptual metrics, fully consistent with our qualitative and statistical observations.
5 Conclusion
In this paper, we propose LPNSR, an efficient prior-enhanced diffusion SR framework. We first establish a unified MLE-based theoretical framework to derive the closed-form optimal intermediate noise for general diffusion models, and instantiate it to the residual-shifting diffusion paradigm with an LR-guided noise predictor and high-quality pre-upsampling initialization. Extensive experiments show that our 4-step LPNSR achieves SOTA perceptual performance on both synthetic and real-world datasets without external text-to-image priors, and supports flexible 1-4 step arbitrary inference. The core optimal noise derivation paradigm can be generalized to other diffusion frameworks, and we leave efficient training schemes for long-trajectory models to future work.
Appendix A Appendix
In the appendix, we provide the following materials:
  • Extension to the DDPM Paradigm
  • Quantitative and qualitative results of different noise injection strategies.
  • Ablation study on the loss function.
  • Different pre-upsampling backbones for the 4-step diffusion SR.
  • More qualitative comparisons with state-of-the-art methods.
  • The complete training and inference algorithms of our LPNSR framework.
A.1 Extension of optimal intermediate noise derivation to the DDPM paradigm
In this section, we extend the unified maximum likelihood estimation (MLE)-based optimal intermediate noise derivation framework from the main paper to the DDPM paradigm (Ho et al., 2020), following the unified notation system for conditional diffusion models established in the main paper. For the conditional DDPM framework, we follow its native notation: let $\beta_{t}\in(0,1)$ be the predefined noise schedule, with $\alpha_{t}=1-\beta_{t}$ and $\bar{\alpha}_{t}=\prod_{s=1}^{t}\alpha_{s}$ as the cumulative product of $\alpha_{t}$. The forward diffusion process corrupts $x_{0}$ with Gaussian noise, with a closed-form marginal distribution at timestep $t$:
$q(x_{t}|x_{0},y_{0})=\mathcal{N}\left(x_{t};\sqrt{\bar{\alpha}_{t}}x_{0},(1-\bar{\alpha}_{t})\cdot\boldsymbol{I}\right),$ (16)
with corresponding forward marginal mean $\mu_{t}^{m}(x_{0},y_{0})=\sqrt{\bar{\alpha}_{t}}x_{0}$ and marginal variance $\sum_{t}^{m}=1-\bar{\alpha}_{t}$. For the DDPM reverse denoising process, the single-step transition follows an isotropic Gaussian distribution, with the reverse mean parameterized by the denoising network $\epsilon_{\theta}$:
$\mu_{t}^{\theta}(x_{t},y_{0},t)=\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},y_{0},t)\right).$ (17)
The reverse variance $\sigma_{t}^{2}$ follows the original DDPM fixed setting, corresponding to $\sqrt{\sum_{t}}=\sigma_{t}$ in our unified notation, with the reverse iteration given by $x_{t-1}=\mu_{t}^{\theta}(x_{t},y_{0},t)+\sigma_{t}\cdot z_{t-1}$, where $z_{t-1}$ is the intermediate noise universally sampled from a standard Gaussian distribution in conventional DDPM pipelines. Substituting the DDPM formulation into our general optimal noise solution in Eq. (9), we directly obtain the closed-form optimal intermediate noise for the DDPM paradigm:
$z_{t-1}^{*}=\frac{\sqrt{\bar{\alpha}_{t-1}}x_{0}-\frac{1}{\sqrt{\alpha_{t}}}\left(x_{t}-\frac{1-\alpha_{t}}{\sqrt{1-\bar{\alpha}_{t}}}\epsilon_{\theta}(x_{t},y_{0},t)\right)}{\sigma_{t}}.$ (18)
This result aligns perfectly with the core conclusions of the main paper, proving that the optimal intermediate noise for DDPM is a deterministic mapping rather than unconstrained random Gaussian noise, verifying that our MLE-based derivation framework is generalizable across mainstream diffusion paradigms beyond the residual-shifting framework, and provides theoretical guidance for optimal noise design in long-trajectory DDPM-based SR models.
A.2 Validation of SR-Based approximate optimal noise
In this section, we verify the feasibility of generating approximate optimal noise via a SR image as a proxy for the ground-truth HR image. Specifically, we use the pre-upsampled output of SwinIR-GANLiang et al. (2021) as the $x_{0}$ substitute in Eq. (12) to generate noise, and perform the full 4-step inference to produce the final result. We compare its performance with that of random Gaussian noise, theoretical optimal noise (calculated from the ground-truth HR image), and our LR-guided noise predictor, with results presented in Tab. 5. It can be seen that the theoretical optimal noise achieves perfect pixel-level reconstruction of the ground-truth HR image, and the approximate optimal noise significantly improves reconstruction fidelity, while its perceptual quality is inferior to our trained noise predictor. Fig. 5 presents the qualitative comparison of all noise injection strategies. Our LR-guided noise predictor produces results that closely align with the theoretical optimal noise, faithfully recovering fine details and structures consistent with the LR input. In comparison, the approximate optimal noise shows limited perceptual quality, while unconstrained random Gaussian noise results in severe misalignment between the generated textures and structures and the input LR image.
A.3 Ablation study on the loss functions
Tab. 6 presents the ablation results of our loss function on the ImageNet-Test dataset. The L1 loss alone ensures optimal pixel fidelity but leads to poor perceptual quality; the LPIPS loss balances fidelity and visual similarity, while the GAN loss significantly enhances image realism. Our final combined loss achieves the best trade-off between pixel-level fidelity and perceptual realism, which is the core reason for adopting this configuration in our study.
A.4 Pre-Upsampling backbones
We evaluate the performance of our 4-step diffusion SR framework equipped with different pre-upsampling backbones, with quantitative results presented in Tab. 7. All three tested networks (BSRGANZhang et al. (2021), RealESRGANWang et al. (2021), SwinIR-GANLiang et al. (2021)) deliver comparable fidelity performance on both ImageNet-Test and RealSR datasets, verifying the good compatibility of our framework. Among them, SwinIR-GAN achieves superior perceptual performance on all non-reference metrics across both datasets, while maintaining competitive PSNR and SSIM. This validates the superiority of SwinIR-GAN in balancing fidelity and visual realism for our diffusion SR pipeline, and we thus adopt it as the default pre-upsampling initialization network in our framework.
A.5 More qualitative comparisons
Fig. 6 and Fig. 7 presents more qualitative comparisons of our methods against recent SOTA methods. One can see that our LPNSR achieves comparable or superior visual quality to T2I-utilizing methods such as OSEDiffWu et al. (2024a) and InvSRYue et al. (2025), without relying on any external priors.
A.6 Training and inference algorithms
The pseudo-code of the LPNSR framework training and inference algorithms is summarized in Alg. 1 and 2.
Figure 5: Qualitative comparison of different noise injection strategies. (Zoom in for best view)
Table 5: Quantitative comparison of different intermediate noise injection strategies on ImageNet-Test and RealSR datasets.
[…]
ImageNet-Test Random Gaussian Noise 24.70 0.6567 0.3043 6.8422 4.4138 0.7341 70.8225
Approximate Optimal Noise 27.46 0.7636 0.2242 5.3760 4.3798 0.5679 66.2120
LR-Guided Noise Predictor 26.11 0.7054 0.2424 4.3807 3.1995 0.6921 71.7105
Theoretical Optimal Noise 34.61 0.9282 0.0452 5.1169 4.1068 0.5623 65.3368
RealSR Random Gaussian Noise 22.68 0.6194 0.4160 6.9618 5.1249 0.7162 60.9120
Approximate Optimal Noise 27.02 0.7942 0.2597 5.4754 5.3232 0.4993 60.0818
LR-Guided Noise Predictor 24.62 0.7003 0.3229 4.2175 3.6963 0.7180 67.5634
Theoretical Optimal Noise 35.83 0.9723 0.0338 6.0362 5.5348 0.4684 58.6619
Table 6: Quantitative ablation studies on the loss function, wherein the hyper-parameters $\lambda_{l}$ and $\lambda_{g}$ control the weight importance of the LPIPS loss and the GAN loss, respectively. The results are evaluated on the ImageNet-Test dataset under the 4-step sampling setting.
Methods Hyper-parameters Metrics
[…]
Baseline1 0.0 0.0 27.20 0.7265 0.2823 5.2234 3.8354 0.6268 66.6248
Baseline2 1.0 0.0 26.70 0.7158 0.2643 4.7588 3.5213 0.6621 69.5726
Baseline3 0.0 0.1 25.95 0.7003 0.2513 4.4315 3.2229 0.7044 72.1065
LPNSR 1.0 0.1 26.11 0.7054 0.2424 4.3807 3.1995 0.6921 71.7105
Table 7: Quantitative comparison of different pre-upsampling networks for the 4-step diffusion SR on ImageNet-Test and RealSR.
[…]
ImageNet-Test BSRGANZhang et al. (2021) 26.08 0.7052 0.2439 4.4115 3.2022 0.6837 71.7541
RealESRGANWang et al. (2021) 26.14 0.7066 0.2411 4.4835 3.2214 0.6774 71.4783
SwinIR-GANLiang et al. (2021) 26.11 0.7054 0.2424 4.3807 3.1995 0.6921 71.7105
RealSR BSRGANZhang et al. (2021) 24.66 0.7009 0.3243 4.2239 3.6954 0.7159 67.5517
RealESRGANWang et al. (2021) 24.59 0.7001 0.3256 4.2375 3.7023 0.7123 67.4834
SwinIR-GANLiang et al. (2021) 24.62 0.7003 0.3229 4.2175 3.6963 0.7180 67.5634
Figure 6: More visualization comparisons of different models. (Zoom in for best view)
Figure 7: More visualization comparisons of different models. (Zoom in for best view)
Algorithm 1 Noise Predictor Training
1:HR/LR image pairs $\mathcal{D}$, pretrained UNet denoiser(frozen), optimizer $\mathcal{O}$, loss $\mathcal{L}$, Initialize $g_{w}$, sampling steps $T$
2:Trained noise predictor $g_{w}$
3:while not converged do
4:  Sample $x_{0},y_{0}\sim\mathcal{D}$
5:  Sample $z_{T}\sim\mathcal{N}(0,I)$, $\hat{y_{0}}=\text{Bicubic}(y_{0})$
6:  $x_{T}=\hat{y_{0}}+\kappa\sqrt{\eta_{T}}z_{T}$
7:  for $t=T,T-1,\dots,1$ do
8:   if $t>1$ then
9:     $x_{0}^{\prime}=\text{UNet}(x_{t},y_{0},t)$
10:     $\mu=\frac{\eta_{t-1}}{\eta_{t}}x_{t}+\frac{\alpha_{t}}{\eta_{t}}x_{0}^{\prime}$
11:     $x_{t-1}=\mu+\sqrt{\Sigma_{\theta}}\cdot g_{w}(x_{t},x_{0}^{\prime},y_{0},t)$
12:   else
13:     $x_{0}^{\prime}=\text{UNet}(x_{t},y_{0},t)$
14:   end if
15:  end for
16:  Compute loss $\mathcal{L}(x_{0}^{\prime},x_{0})$, $\mathcal{O}.\text{step}(\mathcal{L})$
17:end while
18:return $g_{w}$
Algorithm 2 Inference
1:LR image $y_{0}$, pretrained UNet denoiser, noise predictor $g_{w}$, pretrained SR network, sampling steps $T$
2:Generated HR image $x_{0}^{\prime}$
3:Sample $z_{T}\sim\mathcal{N}(0,I)$, $\hat{y_{0}}=\text{pre-upsample}(y_{0})$
[…]
5:for $t=T,T-1,\dots,1$ do
6:  if $t>1$ then
7:   $x_{0}^{\prime}=\text{UNet}(x_{t},y_{0},t)$
8:   $\mu=\frac{\eta_{t-1}}{\eta_{t}}x_{t}+\frac{\alpha_{t}}{\eta_{t}}x_{0}^{\prime}$
9:   $x_{t-1}=\mu+\sqrt{\Sigma_{\theta}}\cdot g_{w}(x_{t},x_{0}^{\prime},y_{0},t)$
10:  else
11:   $x_{0}^{\prime}=\text{UNet}(x_{t},y_{0},t)$
12:  end if
13:end for
14:return $x_{0}^{\prime}$